Silver Company makes a product that is very popular as a Mother’s Day gift. Thus, peak sales occur in May of each year, as shown in the company’s sales budget for the second quarter given below:

April May June Total
Budgeted sales (all on account) $380,000 $580,000 $200,000 $1,160,000

From past experience, the company has learned that 20% of a month’s sales are collected in the month of sale, another 60% are collected in the month following sale, and the remaining 20% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $310,000, and March sales totaled $340,000.

Required:

1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

2. What is the accounts receivable balance on June 30th?

evue balance O ule Sou Answer is not complete. Complete this question by entering your answers in the tabs below. Required 1 Required 2 Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter. Schedule of Expected Cash Collections April May June Total $ 62,000 204,000 68,000 February sales March sales April sales May sales June sales Total cash collections 62,000 272,000 76,000228,00076,000380,000 348000 348,000 $ 342,000 296,000 $ 424,000 $ 1,062,000 Required 1 Required 2> くPrev 1016 İİ, Next>

From past experience, the company has learned that 20% of a months sales are collected in the collected in the month following sale, and the remaining 20% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $310,000, and March sales totaled $340,000. month of sale, another 60% ar Required: 1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter 2. What is the accounts receivable balance on June 30th? Answer is not complete. Complete this question by entering your answers in the tabs below. Required 1 Required 2 What is the accounts receivable balance on June 30th? $276,000 < Required 1 Required 2 < Prev 1of 6 ::: Next >
Answer #1
Schedule of expected cash collections
April May June Total
February sales $            62,000 $       62,000
March sales $         2,04,000 $            68,000 $   2,72,000
April sales $            76,000 $         2,28,000 $         76,000 $   3,80,000
May sales $         1,16,000 $     3,48,000 $   4,64,000
June sales $         40,000 $       40,000
Total cash collections $         3,42,000 $         4,12,000 $     4,64,000 $ 12,18,000
Total Accounts receivables at June 30 = $276000
Workings:
Schedule of expected cash collections
April May June Total
February sales =310000*0.2
March sales =340000*0.6 =340000*0.2
April sales =380000*0.2 =380000*0.6 =380000*0.2
May sales =580000*0.2 =580000*0.6
June sales =200000*0.2
Total Accounts receivables at June 30 = ($580000*0.2) + ($200000*0.8) = $276000